# Fine-Grained Image-Text Alignment in Medical Imaging Enables Cyclic Image-Report Generation

### v4

@@ Abstract
Fine-grained vision-language models (VLM) have been widely used for inter-modality local alignment between the predefined fixed patches and textual words. However, in the medical analysis, lesions exhibit varying sizes and positions, and the fixed patches may cause incomplete representations of lesions. Moreover, these methods provide explainability by using heatmaps to show the general potential image areas associated with texts rather than specific regions, making their explanations not explicit and specific enough. To address these issues, we propose a novel Adaptive patch-word Matching (AdaMatch) model to correlate chest X-ray (CXR) image regions with words in medical reports and apply it to CXR-report generation to provide explainability for the generation process. AdaMatch exploits the fine-grained relation between adaptive patches and words to provide explanations of specific image regions with corresponding words. To capture the abnormal regions of varying sizes and positions, we introduce the Adaptive Patch extraction (AdaPatch) module to acquire the adaptive patches for these regions adaptively. In order to provide explicit explainability for CXR-report generation task, we propose an AdaMatch-based bidirectional large language model for Cyclic CXR-report generation (AdaMatch-Cyclic). It employs the AdaMatch to obtain the keywords for CXR images and ‘keypatches’ for medical reports as hints to guide CXR-report generation. Extensive experiments on two publicly available CXR datasets prove the effectiveness of our method and its superior performance to existing methods.
@@ 1 Introduction
Inter-modality alignment, such as vision and language, has been an important task with growing interests in the field of computer vision, especially with the recent advancement in representation learning [34]. Technologies like contrastive learning and self-supervised learning have dramatically improved state-of-the-art alignment performance. Recent vision-language models (VLMs) demonstrate two approaches: global contrastive alignment, which integrates images and texts at a global level [34, 19, 17, 44, 46], and local alignment, focusing on detailed connections between visual objects and textual words [6, 25, 24, 52, 21, 50], as illustrated in Fig. 1.
@@
Figure 1: Current vision-language models (VLM) achieve (a) global alignment and (b) local alignment by matching overall visual with textual features, and aligning patch with word features, respectively. (c) To exploit the relation between textual words and abnormal patches with different sizes, the proposed AdaMatch obtains adaptive patch features and aligns them with word features.
@@
Current VLMs with local alignment either adopt the pre-trained object detector to extract region-of-interest (ROI) features from images and match the corresponding object features with textual words [6, 25, 24, 52]; or align the visual token from each patch and the textual token into the same embedding space [21, 50, 18, 41]. The former highly relies on the quality of the object detector and its predefined classes, which is less generalizable to new domains. The latter family of methods learns the alignment in a more automatic and data-driven manner. However, most of these methods depend on pre-defined patch size and position (e.g., grids) across images. In most challenging cases, such as the analysis of medical image analysis, lesions can exhibit a wide range of shapes, sizes, and positions. A fixed definition of image patches can lead to incomplete or ambiguous representations of the key imaging abnormalities. Hence, it is highly needed to adaptively exploit the fine-grained relationship between image embeddings derived from a more flexible patching scheme and textual embeddings.
@@
Another challenge in the current local medical VLMs lies in their explainability: it is generally difficult to delineate the image-text relationship learned by the model, especially for the current medical VLMs. Current solutions to provide such explanations in medical VLMs leverage the attention maps from the intermediate layer to visualize the location of the abnormalities [32, 16, 45]. Other methods  [40, 5, 28] utilize network gradients such as Grad-CAM [38] to generate the heatmaps according to lesion types based on ground-truth reports. However, both maps can only show the general potential areas associated with the corresponding text data rather than clearly pinpointing a specific region. In addition, gradient-based methods need ground-truth reports, prohibiting them from functioning correctly beyond training data. It is thus highly necessary to develop a mechanism that could provide explicit and specific explanations of the input image or text during inference time.
@@
To address these two challenges above, we propose a novel Adaptive patch-word Matching (AdaMatch) model to match fine-grained image regions of various sizes and positions with textual data. AdaMatch introduces an image encoder with multiple Adaptive Patch extraction (AdaPatch) modules to adaptively acquire the patches associated with certain text tokens. It then performs patch-word alignment based on contrastive learning. AdaMatch is specifically developed in the context of aligning radiology images (chest X-ray, CXR) and their corresponding radiology reports with the capability of achieving cyclic (CXR-to-report and report-to-CXR) generation based on the learned alignment. Our premise is that such a cyclic generation task would serve as the best use case and evaluation criterion for the desired fine-grained alignment. Also, the fine-grained cyclic generation between CXR and report will provide a natural explainability for how the model aligns the two modalities: for any given text token, we can visualize its matching imaging manifestation; and for any image region within a CXR image, we can tell the type of lesion or anatomical region it belongs to.
@@
To implement the cyclic CXR-report generation, we propose an AdaMatch-based bidirectional model (AdaMatch-Cyclic). AdaMatch-Cyclic employs AdaMatch to identify the keywords for CXR images and ‘keypatches’ for medical reports to guide the generation tasks. Since the potential keywords for CXR images cover a wide range and ground-truth reports cannot be used during inference, we predefine a textual codebook with the most common entities from medical reports as prior knowledge during fine-grained alignment. With the textual codebook, AdaMatch aligns it with the adaptive patches to obtain matched keywords as guidance for report generation. Next, a VQ-GAN model encodes the CXR image into image tokens, and a Large Language Model (LLM) takes image tokens, the matched keywords, and the instructions as input to generate the medical reports. Similarly, we also build a visual codebook with the most commonly seen patches as ‘keypatches’, and use AdaMatch to obtain the matched keypatches from given text reports as hints for CXR generation. With medical reports, matched keypatches, and instructions, LLM outputs the image tokens, and then the VQ-GAN model decodes them into the generated CXR image. Our contributions are summarized as follows:
@@
  • To exploit the fine-grained relation between the CXR image patches with words of medical reports, we propose an Adaptive patch-word Matching (AdaMatch) model to obtain adaptive patches for abnormal regions and perform alignment between them and texts in medical reports.
@@
  • We devise an AdaMatch-based bidirectional LLM for Cyclic CXR-report generation (AdaMatch-Cyclic) to facilitate the generation between CXR and reports. Moreover, we build the textual and visual codebook to utilize AdaMatch to extract useful keywords and keypatches for the report and CXR generation, respectively.
@@
  • Experiments on two publicly available chest X-ray datasets demonstrate the effectiveness of our method and its superior performance over state-of-art methods.
@@ 2.1 Fine-Grained Vision-Language Models
Recently, several fine-grained vision-language models (VLM) [6, 25, 24, 52, 21, 50, 18, 41] achieve the local alignment by exploiting the fine-grained relation between visual objects and textual words. Some methods [6, 25, 24, 52] employ the pre-trained object detector to obtain region-of-interest (ROI) features from images and align the object features with textual features, and others [21, 50, 41, 18] aim to locally align the fixed patches with the textual words. For instance, Li et al. [24] utilizes Faster-RCNN [36] to detect visual objects and match ROI embeddings with textual embeddings. Moreover, FILIP [50] achieves cross-modal interaction by maximizing the similarity between the visual tokens of fixed patches and textual tokens. The former requires a precise object detector, and the latter primarily learns the relation between fixed patches inside the predefined grid and words. Nevertheless, when lung lesions differ in size, shape, and position, these methods may split them into separate patches, leading to incomplete semantic information of patches. Thus, we devise an Adaptive patch-word Matching (AdaMatch) model to adaptively exploit the fine-grained relation between flexible patches and textual words.
@@ 2.2 CXR-Report Generation
Medical VLMs are widely used in downstream tasks for chest radiographs, including CXR-to-report generation [7, 8, 48, 42, 39, 47] and report-to-CXR generation [37, 3, 2, 22, 23] tasks. CXR-to-report generation has been investigated to reduce the burden of radiologists and promote clinical automation. For instance, Chen et al. [8] introduces a cross-modal memory network with shared memory to align images with texts to promote report generation performance. In report-to-CXR generation task, previous methods generate high-quality and annotated CXR images from medical reports, to provide more training data and avoid privacy laws and considerations. Some diffusion-based methods [2, 22] take the impression section of reports as input, ignoring the finding section that includes detailed descriptions. To consider more details in reports, several transformer-based methods [22, 23] take both finding and impression sections as input to synthesize CXR. In this paper, we use the cyclic generation (i.e. CXR-to-report and report-to-CXR generation) as use case and evaluation criteria for our fine-grained alignment method (AdaMatch). We also design an AdaMatch-Cyclic to employ AdaMatch to improve the explainability of cyclic generation.
@@ 3 Methods
In Fig. 2, we propose an adaptive patch-word matching (AdaMatch) model to associate the CXR image regions with words in medical reports, and apply it to CXR-report generation to provide explainability. Given an input CXR image $x^{I}$, the image encoder with several Adaptive Patch extraction modules (AdaPatch) predicts the location and scale of multiple adaptive patches and processes the adaptive patch embeddings $z^{I}$. With $z^{I}$ and text embeddings $z^{T}$ extracted by text encoder for the medical report $x^{T}$, we compute the similarities $s^{I}$, $s^{T}$ between adaptive patches and text tokens to calculate the contrastive loss $L_{I2T},L_{T2I}$ to optimize AdaMatch model. For CXR-to-report generation, we utilize the frozen AdaMatch model to match a predefined textual codebook with the input CXR image $x^{I}$ to obtain matched keywords $k^{T}$ and feed a LLM with $k^{T}$, the instruction, and image tokens encoded by VQ-GAN from $x^{I}$ to generate a medical report $x^{T^{\prime}}$. Then, the AdaMatch model is used to match a predefined visual codebook with the generated report $x^{T^{\prime}}$ to acquire the matched keypatches $k^{I}$. With instruction, $x^{T^{\prime}}$ and $k^{I}$, LLM outputs image tokens of the generated CXR image $x^{I^{\prime}}$, which is optimized by $L_{text}$ and $L_{img}$.
@@ 3.1 Adaptive Patch-Word Matching (AdaMatch)
Current VLMs [6, 25, 24, 52, 21, 50] achieve fine-grained alignment between visual objects and textual words, but they may split lung lesions into separate fixed patches due to various shapes, sizes, and positions of lung lesions. Thus, we propose an Adaptive patch-word Matching (AdaMatch) to automatically locate the important regions, extract adaptive patches for these regions, and align them with corresponding textual words. As shown in Fig. 2 (a), AdaMatch includes an image encoder to extract adaptive patch features and a text encoder to encode textual features.
Image Encoder. To obtain the adaptive patches, the image encoder consists of four stages with feature maps of decreasing scales. Specifically, in the first stage, we first adopt the patch embedding module to split the CXR image $x^{I}\in\mathbb{R}^{H\times W\times C}$ into $N$ patches with fixed size $s\times s$ and project them to patch embeddings $z^{(i)}(1\leq i\leq N)$ through a fully connected layer $g$, $z^{(i)}=g([a^{(i,1)};\dots;a^{(i,s\times s)}])$, where $a^{(i,j)}$ indicates the image features for the pixel located at $q^{(i,j)}$ and []delimited-[][$\cdot$][ ⋅ ] denotes the concatenation operation. Afterward, we pass the patch embeddings $z^{(i)}$ with the position embeddings $e_{pos}$ into a transformer encoder [43] $R$ to obtain the outputs, $z^{(i)}=R(z^{(i)}+e_{pos})$, where $R$ includes an attention layer and a feed-forward layer. In order to locate the potential lung lesions, we devise an Adaptive Patch Extraction module (AdaPatch) for the rest of stages. In AdaPatch, a fully connected layer is fed with patch embeddings $z^{(i)}$, and then the outputs are passed through four separate fully connected layers $f_{1},f_{2},f_{3},f_{4}$ to predict the offset $(\delta_{x}^{(i)},\delta_{y}^{(i)})$ and patch size $(s_{w}^{(i)},s_{h}^{(i)})$ for the adaptive patches, where the offset indicates the shift to the center of original fixed patches $(c^{(i)}_{x},c^{(i)}_{y})$,
$\delta_{x}^{(i)}=\small{Tanh}(f_{1}(z^{(i)})),\delta_{y}^{(i)}=\small{Tanh}(f_{2}(z^{(i)})),$ (1)
$s_{w}^{(i)}=\small{ReLU}(\small{Tanh}(f_{3}(z^{(i)})),s_{h}^{(i)}=\small{ReLU}(\small{Tanh}(f_{4}(z^{(i)})).$ (2)
@@
and uniformly sample $m\times m$ feature points inside the patches. Since the coordinates may be fractional, bilinear interpolation is used to obtain the sampled feature points $\left\{\hat{p}^{(j)}\right\}_{1\leq j\leq m\times m}$. The embeddings of all sampled points are flattened and fed into a fully connected layer $f_{5}$ to obtain patch embeddings, $z^{(i)}=f_{5}([\hat{p}^{(i,1)};\dots;\hat{p}^{(i,m\times m)}])$. Finally, we pass $z^{(i)}$ with position embeddings $e_{pos}$ into a transformer encoder $R$. The final adaptive patch embeddings $z^{I}\in\mathbb{R}^{N\times d}$ are the ensemble of $z^{(i)}(1\leq i\leq N)$, $z^{I}=E^{I}(x^{I})=\left\{z^{I}_{1},\cdots,z^{I}_{N}\right\},$ where $E^{I}$, $x^{I}$, and $d$ denote the image encoder, the input CXR image, and the dimension of adaptive patch embeddings, respectively.
Text Encoder. We adopt the BioClinicalBERT [1] as the text encoder $E^{T}$ to encode the medical report $x^{T}$ into text embeddings $z^{T}$. To make $z^{T}$ with the same dimension as $z^{I}$, we feed a fully connected layer with $z^{T}$ to reduce its dimension, $z^{T}=E^{T}(x^{T})=\left\{z^{T}_{1},\cdots,z^{T}_{K}\right\}(z^{T}\in\mathbb{R}^{K\times d})$, where $K$ denotes the number of text tokens.
Patch-Word Alignment. In order to exploit the relation between the adaptive patches and textual tokens, we perform fine-grained contrastive representation learning to achieve patch-word alignment. Concretely, for the $i$-th CXR image $x^{I}_{i}$ and $j$-th medical report $x^{T}_{j}$, we first compute the similarities between all the adaptive patch embeddings $z^{I}_{n}(1\leq n\leq N)$ and all the text embeddings $z^{T}_{k}(1\leq k\leq K)$, and use the largest similarity $\max_{(1\leq k\leq K)}{{(z^{I}_{n})}^{\top}z^{T}_{k}}$ as the patch-word maximum similarity for $n$-th adaptive patch embedding. Then, the patch-word maximum similarities for all the adaptive patch embeddings are averaged as the similarity $s^{I}_{i,j}$ of the $i$-th CXR image to the $j$-th medical report,
$s^{I}_{i,j}=\frac{1}{N}\displaystyle\sum_{n=1}^{N}(z^{I}_{n})^{\top}z^{T}_{m_{n}^{I}},$ (5)
where $m_{n}^{I}=\arg\max_{(1\leq k\leq K)}{{(z^{I}_{n})}^{\top}z^{T}_{k}}$. Similarly, the similarity of the $j$-th medical report to the $i$-th CXR image is defined as,
$s^{T}_{i,j}=\frac{1}{K}\displaystyle\sum_{k=1}^{K}(z^{I}_{m_{k}^{T}})^{\top}z^{T}_{k},$ (6)
where $m_{k}^{T}=\arg\max_{(1\leq n\leq N)}{{(z^{I}_{n})}^{\top}z^{T}_{k}}$. We exclude the padded textual tokens when computing the similarity.
@@
where $b$ denotes the batch size, $\tau$ represents temperature hyperparameter and $\left\{x_{i}^{I},x_{i}^{T}\right\}$ indicates the positive CXR-report pair. Similarly, the adaptive word-patch contrastive loss $L_{i}^{T}$ for $i$-th medical report $x_{i}^{T}$ is formulated as,
@@
The final contrastive loss for a mini-batch is calculated by,
$L=\frac{1}{2}\sum_{i=1}^{b}{(L_{i}^{I2T}+L_{i}^{T2I})}.$ (9)
With the final contrastive loss, the image and text encoders are optimized to automatically locate the important patches with varied sizes and exploit the fine-grained relation between the patches and words.

### v5

# Fine-Grained Image-Text Alignment in Medical Imaging Enables Explainable Cyclic Image-Report Generation

@@ Abstract
Fine-grained vision-language models (VLM) have been widely used for inter-modality local alignment between the predefined fixed patches and textual words. However, in medical analysis, lesions exhibit varying sizes and positions, and using fixed patches may cause incomplete representations of lesions. Moreover, these methods provide explainability by using heatmaps to show the general image areas potentially associated with texts rather than specific regions, making their explanations not explicit and specific enough. To address these issues, we propose a novel Adaptive patch-word Matching (AdaMatch) model to correlate chest X-ray (CXR) image regions with words in medical reports and apply it to CXR-report generation to provide explainability for the generation process. AdaMatch exploits the fine-grained relation between adaptive patches and words to provide explanations of specific image regions with corresponding words. To capture the abnormal regions of varying sizes and positions, we introduce an Adaptive Patch extraction (AdaPatch) module to acquire adaptive patches for these regions adaptively. Aiming to provide explicit explainability for the CXR-report generation task, we propose an AdaMatch-based bidirectional LLM for Cyclic CXR-report generation (AdaMatch-Cyclic). It employs AdaMatch to obtain the keywords for CXR images and ‘keypatches’ for medical reports as hints to guide CXR-report generation. Extensive experiments on two publicly available CXR datasets validate the effectiveness of our method and its superior performance over existing methods.
Fine-Grained Image-Text Alignment in Medical Imaging
Enables Explainable Cyclic Image-Report Generation
Wenting Chen1  Linlin Shen3  Jingyang Lin4  Jiebo Luo4 Xiang Li5  Yixuan Yuan2thanks: Xiang Li and Yixuan Yuan are corresponding authors. 1City University of Hong Kong   2The Chinese University of Hong Kong 3Shenzhen University   4University of Rochester 5Massachusetts General Hospital and Harvard Medical School 1wentichen7-c@my.cityu.edu.hk  2yxyuan@ee.cuhk.edu.hk  3llshen@szu.edu.cn 4{jluo@cs, jlin81@ur}.rochester.edu 5xli60@mgh.harvard.edu
@@ 1 Introduction
Inter-modality alignment, such as vision and language, has been an important task with growing interests in the field of computer vision, especially with the recent advancement in representation learning Radford et al. (2021). Technologies like contrastive learning and self-supervised learning have dramatically improved state-of-the-art alignment performance. Recent vision-language models (VLMs) demonstrate two approaches: global contrastive alignment, which integrates images and texts at a global level Radford et al. (2021); Jia et al. (2021); Jang et al. (2023); Wang et al. (2023); Yang et al. (2022), and local alignment, focusing on detailed connections between visual objects and textual words Chen et al. (2020a); Li et al. (2020b, a); Zhan et al. (2021); Kim et al. (2021); Yao et al. (2021), as illustrated in Fig. 1.
@@
Figure 1: Current vision-language models (VLM) achieve (a) global alignment and (b) local alignment by matching overall visual with textual features, and aligning patches with word features, respectively. (c) To exploit the relation between textual words and abnormal patches with varied sizes, our AdaMatch obtains adaptive patch features and aligns them with word features.
@@
Current VLMs with local alignment either adopt the pre-trained object detector to extract region-of-interest (ROI) features from images and match the corresponding object features with textual words Chen et al. (2020a); Li et al. (2020b, a); Zhan et al. (2021), or align the visual token from each patch and the textual token into the same embedding space Kim et al. (2021); Yao et al. (2021); Ji et al. (2021); Wang et al. (2022a). The former highly relies on the quality of the object detector and its predefined classes, which is less generalizable to new domains. The latter family of methods learns the alignment in a more automatic and data-driven manner. However, most of these methods depend on a pre-defined patch size and positions (e.g., grids) across images. In the most challenging cases, such as the analysis of medical image, lesions can exhibit a wide range of shapes, sizes, and positions. A fixed partition of image patches can lead to incomplete or ambiguous representations of the key imaging abnormalities. Therefore, it is highly desirable to adaptively exploit the fine-grained relationship between image embeddings derived from a more flexible patching scheme and textual embeddings.
@@
Another challenge in the current VLMs lies in their explainability: it is generally difficult to delineate the image-text relationship learned by the model, especially for the current medical VLMs. Current solutions to provide such explanations in medical VLMs leverage the attention maps from the intermediate layer to visualize the location of the abnormalities Moon et al. (2022); Huang et al. (2021); Yan and Pei (2022). Other methods  Wan et al. (2023); Chen et al. (2023); Liu et al. (2023) utilize network gradients such as Grad-CAM Selvaraju et al. (2017) to generate the heatmaps according to lesion types based on ground-truth reports. However, both maps can only show the general areas potentially associated with the corresponding text data rather than pinpointing a specific region. In addition, gradient-based methods need ground-truth reports, prohibiting them from functioning correctly beyond training data. It is thus highly necessary to develop a mechanism that could provide explicit and specific explanations of input image or text during inference time.
@@
To address these two challenges above, we propose a novel Adaptive patch-word Matching (AdaMatch) model to match fine-grained image regions of various sizes and positions with textual data. AdaMatch introduces an image encoder with multiple Adaptive Patch extraction (AdaPatch) modules to adaptively acquire the patches associated with certain text tokens. It then performs patch-word alignment based on contrastive learning. AdaMatch is specifically developed in the context of aligning radiology images (chest X-ray, CXR) and their corresponding radiology reports with the capability of achieving cyclic (CXR-to-report and report-to-CXR) generation based on the learned alignment. Our premise is that such a cyclic generation task would serve as the best use case and evaluation criterion for the desired fine-grained alignment. Also, fine-grained cyclic generation between CXR and report will provide natural explainability for how the model aligns two modalities: for any given text token, we can visualize its matching imaging manifestation; and for any image region within a CXR image, we can tell the type of lesion or anatomical region it belongs to.
@@
To implement the cyclic CXR-report generation, we propose an AdaMatch-based bidirectional model (AdaMatch-Cyclic). AdaMatch-Cyclic employs AdaMatch to identify the keywords for CXR images and the ‘keypatches’ for medical reports to guide the generation tasks. Since the potential keywords for CXR images cover a wide range and ground-truth reports cannot be used during inference, we predefine a textual codebook with the most common entities from medical reports as prior knowledge during fine-grained alignment. With the textual codebook, AdaMatch aligns it with the adaptive patches to obtain matched keywords to facilitate report generation. Next, a VQ-GAN model encodes the CXR image into image tokens, and a Large Language Model (LLM) takes image tokens, the matched keywords, and the instructions as input to generate medical reports. Similarly, we also build a visual codebook with the most commonly seen patches as ‘keypatches’, and use AdaMatch to obtain the matched keypatches from given text reports as hints for CXR generation. Utilizing medical reports, matched keypatches, and instructions, LLM generates image tokens, subsequently decoded by the VQ-GAN model to produce the resulting CXR image. Our contributions are summarized as follows:
@@
  • To exploit the fine-grained relation between CXR image patches and words of medical reports, we propose an Adaptive patch-word Matching (AdaMatch) model to obtain adaptive patches for abnormal regions and perform alignment between them and texts in medical reports.
@@
  • We devise an AdaMatch-based bidirectional LLM for Cyclic CXR-report generation (AdaMatch-Cyclic) to facilitate the bi-directional generation between CXR and reports. Moreover, we build the textual and visual codebook to utilize AdaMatch to extract useful keywords and keypatches for the report and CXR generation, respectively.
@@
  • Experiments on two publicly available chest X-ray datasets demonstrate the effectiveness of our method and its superior performance over the state-of-art methods.
@@ 2.1 Fine-Grained Vision-Language Models
Recently, several fine-grained vision-language models (VLM) achieve the local alignment by exploiting the fine-grained relation between visual objects and textual words. Some methods Chen et al. (2020a); Li et al. (2020b, a); Zhan et al. (2021) employ the pre-trained object detector to obtain object features from images and align them with textual features, and others Kim et al. (2021); Yao et al. (2021); Wang et al. (2022a); Ji et al. (2021); Xu et al. (2023); Jiang et al. (2023) aim to align the fixed patches with the textual words locally. The former relies on a precise object detector, and the latter focuses on the relation between fixed patches inside the predefined grid and words. However, varying lung lesion characteristics may cause these methods to divide them into separate patches, resulting in incomplete semantic information. Thus, we devise an Adaptive patch-word Matching (AdaMatch) model to adaptively exploit the fine-grained relation between flexible patches and textual words.
@@ 2.2 CXR-Report Generation
Medical VLMs are widely used in downstream tasks for chest radiographs, including CXR-to-report generation Chen et al. (2020b, 2021a); Yang et al. (2021); Wang et al. (2022b); Voutharoja et al. (2023); Yang et al. (2023); Shi et al. (2023); Huang et al. (2023) and report-to-CXR generation Rombach et al. (2022); Chambon et al. (2022b, a); Lee et al. (2023a, b); Han et al. (2024); Shentu and Al Moubayed (2024); Hou et al. (2023); Hashmi et al. (2024); Chen et al. (2024) tasks. For CXR-to-report generation, Chen et al., 2021a introduces a cross-modal memory network with shared memory to align images with texts to promote report generation performance. In the report-to-CXR generation task, prior techniques create annotated CXR images from medical reports to augment training data and address privacy concerns, which are categorized into diffusion-based and transformer-based methods. In this paper, we use the cyclic generation (i.e. CXR-to-report and report-to-CXR generation) as use case and evaluation criteria for our fine-grained alignment method (AdaMatch). We also design an AdaMatch-Cyclic to employ AdaMatch to improve the explainability of cyclic generation.
@@ 3 Methods
In Fig. 2, we propose an adaptive patch-word matching (AdaMatch) model to associate CXR image regions with words in medical reports, and apply it to CXR-report generation to enhance explainability. Given an input CXR image $x^{I}$, the image encoder with several Adaptive Patch extraction modules (AdaPatch) predicts the location and scale of multiple adaptive patches and processes the adaptive patch embeddings $z^{I}$. With $z^{I}$ and text embeddings $z^{T}$ extracted by text encoder for the medical report $x^{T}$, we compute the similarities $s^{I}$, $s^{T}$ between adaptive patches and text tokens to calculate the contrastive loss $L_{I2T},L_{T2I}$ to optimize AdaMatch. For CXR-to-report generation, we utilize the frozen AdaMatch model to match a predefined textual codebook with the input CXR image $x^{I}$ to obtain matched keywords $k^{T}$ and feed a LLM with $k^{T}$, the instruction, and image tokens encoded by VQ-GAN from $x^{I}$ to generate a medical report $x^{T^{\prime}}$. Then, the AdaMatch model is used to match a predefined visual codebook with the generated report $x^{T^{\prime}}$ to acquire the matched keypatches $k^{I}$. With instruction, $x^{T^{\prime}}$ and $k^{I}$, LLM outputs image tokens of the generated CXR image $x^{I^{\prime}}$, which is optimized through $L_{text}$ and $L_{img}$.
@@ 3.1 Adaptive Patch-Word Matching (AdaMatch)
Current VLMs Chen et al. (2020a); Yao et al. (2021) achieve fine-grained alignment between visual objects and textual words, but they may split lung lesions into separate fixed patches due to various sizes of lung lesions. Thus, we propose an Adaptive patch-word Matching (AdaMatch) to locate the important regions, extract adaptive patches for these regions, and align them with corresponding textual words, as shown in Fig. 2 (a).
Image Encoder. To obtain the adaptive patches, the image encoder comprises four stages with feature maps of decreasing scales. Specifically, in the first stage, we first adopt the patch embedding module to split the CXR image $x^{I}\in\mathbb{R}^{H\times W\times C}$ into $N$ patches with fixed size $s\times s$ and project them to patch embeddings $z^{(i)}(1\leq i\leq N)$ through a fully connected (FC) layer $g$, $z^{(i)}=g([a^{(i,1)};\dots;a^{(i,s\times s)}])$, where $a^{(i,j)}$ indicates the image features for the pixel located at the $q^{(i,j)}$. $q^{(i,j)}$ indicates the coordinates for the image pixel. []delimited-[][$\cdot$][ ⋅ ] represents the concatenation operation among the features. Afterward, we pass the patch embeddings $z^{(i)}$ with the position embeddings $e_{pos}$ into a transformer encoder $R$ to obtain the outputs, $z^{(i)}=R(z^{(i)}+e_{pos})$. To locate the potential lung lesions, we devise an Adaptive Patch Extraction module (AdaPatch) for the rest of stages. In AdaPatch, patch embeddings $z^{(i)}$ are fed into an FC layer, followed by four separate FC layers $f_{1},f_{2},f_{3},f_{4}$ to predict offsets $(\delta_{x}^{(i)},\delta_{y}^{(i)})$ and patch sizes $(s_{w}^{(i)},s_{h}^{(i)})$ for adaptive patches, with offsets indicating shifts to the center $(c^{(i)}_{x},c^{(i)}_{y})$ of fixed patches,
$\delta_{x}^{(i)}={Tanh}(f_{1}(z^{(i)})),\delta_{y}^{(i)}={Tanh}(f_{2}(z^{(i)})),$ (1)
$s_{w}^{(i)}={ReLU}({Tanh}(f_{3}(z^{(i)})),s_{h}^{(i)}={ReLU}({Tanh}(f_{4}(z^{(i)})).$ (2)
@@
and uniformly sample $m\times m$ feature points inside the patches. Since the coordinates may be fractional, bilinear interpolation is used to obtain the sampled feature points $\left\{\hat{p}^{(j)}\right\}_{1\leq j\leq m\times m}$. The embeddings of all sampled points are flattened and fed into an FC layer $f_{5}$ to obtain patch embeddings, $z^{(i)}=f_{5}([\hat{p}^{(i,1)};\dots;\hat{p}^{(i,m\times m)}])$. Finally, we pass $z^{(i)}$ with position embeddings $e_{pos}$ into a transformer encoder $R$. The final adaptive patch embeddings $z^{I}\in\mathbb{R}^{N\times d}$ are the ensemble of $z^{(i)}(1\leq i\leq N)$, $z^{I}=E^{I}(x^{I})=\left\{z^{I}_{1},\cdots,z^{I}_{N}\right\},$ where $E^{I}$, $x^{I}$, and $d$ denote the image encoder, the CXR image, and the dimension of $z^{I}$, respectively.
Text Encoder. We adopt a pre-trained text encoder $E^{T}$ to encode the medical report $x^{T}$ into text embeddings $z^{T}$. To make $z^{T}$ with the same dimension as $z^{I}$, we feed an FC layer with $z^{T}$ to reduce its dimension, $z^{T}=E^{T}(x^{T})=\left\{z^{T}_{1},\cdots,z^{T}_{K}\right\}(z^{T}\in\mathbb{R}^{K\times d})$, where $K$ is the number of text tokens.
Patch-Word Alignment. To exploit the relation between the adaptive patches and textual tokens, we perform fine-grained contrastive representation learning to achieve patch-word alignment. Concretely, for the $i$-th CXR image $x^{I}_{i}$ and $j$-th medical report $x^{T}_{j}$, we first compute the similarities between all the adaptive patch embeddings $z^{I}_{n}(1\leq n\leq N)$ and all the text embeddings $z^{T}_{k}(1\leq k\leq K)$, and use the largest similarity $\max_{(1\leq k\leq K)}{{(z^{I}_{n})}^{\top}z^{T}_{k}}$ as the patch-word maximum similarity for $n$-th adaptive patch embedding. Then, the patch-word maximum similarities for all the adaptive patch embeddings are averaged as the similarity $s^{I}_{i,j}$ of the $i$-th CXR image to the $j$-th medical report, $s^{I}_{i,j}=\frac{1}{N}\displaystyle\sum_{n=1}^{N}(z^{I}_{n})^{\top}z^{T}_{m_{n}^{I}}$, where $m_{n}^{I}=\arg\max_{(1\leq k\leq K)}{{(z^{I}_{n})}^{\top}z^{T}_{k}}$. Similarly, the similarity of the $j$-th medical report to the $i$-th CXR image is defined as, $s^{T}_{i,j}=\frac{1}{K}\displaystyle\sum_{k=1}^{K}(z^{I}_{m_{k}^{T}})^{\top}z^{T}_{k}$, where $m_{k}^{T}=\arg\max_{(1\leq n\leq N)}{{(z^{I}_{n})}^{\top}z^{T}_{k}}$. We exclude the padded textual tokens when computing the similarity.
@@
where $b$, $\tau$, and $\left\{x_{i}^{I},x_{i}^{T}\right\}$ represent the batch size, temperature hyperparameter and the positive CXR-report pair, respectively. Similarly, the adaptive word-patch contrastive loss $L_{i}^{T}$ for $i$-th medical report $x_{i}^{T}$ is formulated as,
@@
The final contrastive loss for a mini-batch is calculated by, $L=\frac{1}{2}\sum_{i=1}^{b}{(L_{i}^{I2T}+L_{i}^{T2I})}$. With $L$, AdaMatch learns to locate important patches with varied sizes and exploits the fine-grained relation.